# Thompson Sampling via Fine-Tuning of LLMs

### v3

@@ Abstract
Bayesian optimization in large unstructured discrete spaces is often hindered by the computational cost of maximizing acquisition functions due to the absence of gradients. We propose a scalable alternative based on Thompson sampling that eliminates the need for acquisition function maximization by directly parameterizing the probability that a candidate yields the maximum reward. Our approach, Thompson Sampling via Fine-Tuning (ToSFiT) leverages the prior knowledge embedded in prompt-conditioned large language models, and incrementally adapts them toward the posterior. Theoretically, we derive a novel regret bound for a variational formulation of Thompson Sampling that matches the strong guarantees of its standard counterpart. Our analysis reveals the critical role of careful adaptation to the posterior probability of maximality—a principle that underpins our ToSFiT algorithm. Empirically, we validate our method on three diverse tasks: FAQ response refinement, thermally stable protein search, and quantum circuit design. Within a collection of methods covering Bayesian optimization, reinforcement learning, and evolutionary search, ToSFiT exhibits both state-of-the-art sample efficiency and computational efficiency.
@@ 1 Introduction
Figure 1: ToSFiT treats candidate generation as Thompson sampling and fine-tunes the LLM toward the posterior PoM.
@@ 7.1 Limitations and Future Work
Apart from the code provided in the supplementary materials, pseudocode for ToSFiT is given in Algorithm 1. The baselines are also implemented in the code in a self-contained manner. The experimental environments for the three tasks considered is also provided in the code. Moreover, the provided README.md gives a step-by-step tutorial to reproduce all results in the paper. The hardware and software for running the experiments is described in detail in Section B of the Appendix. Furthermore, Section B.3 details the hyperparameter sweep conducted during development of this paper. In the evaluation, statistical significance is indicated via mean and standard error under $25$ seeds, set to $0,\ldots,24$. All theoretical statements are self-contained and include the complete set of assumptions. Both detailed proofs as well as proof sketches for the main results are provided in Section E of the Appendix. We distinguish between novel claims and prior statements by citing the appropriate works.
@@ Acknowledgment
This work is supported by the Swiss National Science foundation (SNF), grant 10002666.

### v4

@@ Abstract
Bayesian optimization in large unstructured discrete spaces is often hindered by the computational cost of maximizing acquisition functions due to the absence of gradients. We propose a scalable alternative based on Thompson sampling that eliminates the need for acquisition function maximization by directly parameterizing the probability that a candidate yields the maximum reward. Our approach, Thompson Sampling via Fine-Tuning (ToSFiT), leverages the prior knowledge embedded in prompt-conditioned large language models, and incrementally adapts them toward the posterior. Theoretically, we derive a novel regret bound for a variational formulation of Thompson Sampling that matches the strong guarantees of its standard counterpart. Our analysis reveals the critical role of careful adaptation to the posterior probability of maximality—a principle that underpins our ToSFiT algorithm. Empirically, we validate our method on three diverse tasks: FAQ response refinement, thermally stable protein search, and quantum circuit design. Within a collection of methods covering in-context Bayesian optimization, reinforcement learning, and evolutionary search, ToSFiT exhibits both state-of-the-art sample efficiency and computational efficiency.  [^1] Code at https://github.com/IBM/thompson-sampling-via-fine-tuning-of-llms.
@@ 1 Introduction
Figure 1: ToSFiT treats candidate generation as Thompson sampling and fine-tunes the LLM to the posterior PoM.
@@ 7.1 Limitations and Future Work
Apart from the complete experimental code base provided publicly at https://github.com/IBM/thompson-sampling-via-fine-tuning-of-llms, pseudocode for ToSFiT is given in Algorithm 1. The baselines are also implemented in the code in a self-contained manner. The experimental environments for the three tasks considered is also provided in the code. Moreover, the provided README.md gives a step-by-step tutorial to reproduce all results in the paper. The hardware and software for running the experiments is described in detail in Section B of the Appendix. Furthermore, Section B.3 details the hyperparameter sweep conducted during development of this paper. In the evaluation, statistical significance is indicated via mean and standard error under $25$ seeds, set to $0,\ldots,24$. All theoretical statements are self-contained and include the complete set of assumptions. Both detailed proofs as well as proof sketches for the main results are provided in Section E of the Appendix. We distinguish between novel claims and prior statements by citing the appropriate works.
@@ Acknowledgment
This work is supported by the Swiss National Science Foundation (SNSF), grant 10002666.